OpenStudy (anonymous):

help plz ! will fan+ medal 。◕ ‿ ◕。 what kind of curves , i can solve there length using

OpenStudy (anonymous):

\(\color{blue}{\Huge L:=\int\limits_{a}^{b} r'(t)^2 dt}\)

OpenStudy (rational):

arc length formula doesnt look right

OpenStudy (rational):

\(L = \int \limits_a^b | r'(t)| dt\)

OpenStudy (rational):

\(L = \int \limits_a^b | r'(t)| dt = \int \limits_a^b \sqrt{x(t)^2 + y(t)^2 + z(t)^2 } dt\)

OpenStudy (rational):

this works for any kind of curves that are continuuous and differentiable

OpenStudy (anonymous):

you are right , it should be \(\color{blue}{\Huge L:=\int\limits_{a}^{b} \sqrt {r'(t).r'(t)} dt}\)

OpenStudy (rational):

sometime evaluating integral might not be possible, in which case you need to approximate... but still you would use the same formula

OpenStudy (anonymous):

but i tried for more than curve it dint work unless for circle and simple lines otherwise evaluating integral looks not possible

OpenStudy (anonymous):

so, its in general but

OpenStudy (anonymous):

not each time i reach the exact right ?

OpenStudy (anonymous):

well it make sense ! so there is specifies type of curves i can get an exact solution ?!

OpenStudy (anonymous):

well thx @rational for brainstorming !

OpenStudy (rational):

we cannot integrate every curve by hand

OpenStudy (rational):

that should be understood :) we can differentiate anything. but we cannot integrate everything

OpenStudy (anonymous):

ok ಠ_ಠ

OpenStudy (rational):

but still the given formula wont fail, u wil use the same formula... but use computation to approximate the integral instead

OpenStudy (anonymous):

it well need more works lol